OpenStudy (anonymous):

please show steps? -2+p/7=-4

OpenStudy (anonymous):

-2 + p/7 = -4 add 2 to each side +2 +2 p/7 = -2 then multiply each side by 7 p = -2 * 7

OpenStudy (anonymous):

thank you!!